two capacitor of problem 2.12 are now connected in series and to a 9.00-V battery. find (a) the value of equivalent capacitance of the combination, (b) the voltage across each capacitor and (c) the charge in each capacitor


1
Expert's answer
2021-06-30T10:19:24-0400

a)

"C=\frac{C_1C_2}{C_1+C_2}=\frac{7\cdot 13}{7+13}\cdot 10^{-6}=4.6~\mu F,"

b)

"U_1=\frac{UC_2}{C_1+C_2}=\frac{9\cdot 13}{7+13}=5.9~V,"

"U_2=\frac{UC_1}{C_1+C_2}=\frac{9\cdot 7}{7+13}=3 .2~V,"

c)

"q_1=q_2=CU=4.6\cdot 10^{-6}\cdot 9=41~\mu C."